# RealitySummary: Exploring On-Demand Mixed Reality Text Summarization and Question Answering using Large Language Models

### v2

@@ Abstract.
Large Language Models (LLMs) are gaining popularity as tools for reading and summarization aids. However, little is known about their potential benefits when integrated with mixed reality (MR) interfaces to support everyday reading assistants. We developed RealitySummary, an MR reading assistant that seamlessly integrates LLMs with always-on camera access, OCR-based text extraction, and augmented spatial and visual responses in MR interfaces. Developed iteratively, RealitySummary evolved across three versions, each shaped by user feedback and reflective analysis: 1) a preliminary user study to understand user perceptions (N=12), 2) an in-the-wild deployment to explore real-world usage (N=11), and 3) a diary study to capture insights from real-world work contexts (N=5). Our findings highlight the unique advantages of combining AI and MR, including an always-on implicit assistant, minimal context switching, and spatial affordances, demonstrating significant potential for future LLM-MR interfaces beyond traditional screen-based interactions.
@@
journalyear: 2024copyright: acmlicensedconference: arXiv; September, 2024; $c$s: Human-centered computing Mixed / augmented reality
Figure 1. We developed RealitySummary, an on-demand mixed reality reading assistant, over three iterations and conducted three different studies. Through this evolution, we highlight three key findings: 1) unique benefits of MR and LLMs, 2) emergent use cases of MR reading assistants, and 3) design implications for mixed reality reading assistants.
@@ 1.1. Reading Augmentation
Human-computer interaction researchers have explored ways to augment physical documents with AR-enhanced content. Examples include HoloDoc (Li et al., 2019), Replicate and Resuse (Gupta et al., 2020), and PaperTrail (Rajaram and Nebeling, 2022), all of which laid the groundwork for augmenting physical paper by superimposing dynamic virtual content onto printed documents. Prior works have demonstrated these concepts on different platforms. For example, Pacer (Liao et al., 2010), MagPad (Xu et al., 2015), Tecahable Reality (Monteiro et al., 2023), and Dually Noted (Qian et al., 2022) use mobile AR for text augmentation, whereas Replicate and Resuse (Gupta et al., 2020) and Wiki-TUI (Wu et al., 2007) leverage head-mounted displays to overlay information. Other examples such as DigitalDesk (Wellner, 1991), DocuDesk (Everitt et al., 2008), Affinity Lens (Subramonyam et al., 2019), Matulic et al. (Matulic and Norrie, 2013; Matulic et al., 2013), and QOOK (Zhao et al., 2014) use projection mapping for document augmentation. At their core, such augmented reading research (Kirner et al., 2012; Billinghurst et al., 2001; Hidayat et al., 2020; Grasset et al., 2008) aims to improve cognitive load (Cheng, 2017), knowledge accumulation (Dünser et al., 2012; Zhang et al., 2019), and spatial visualization (Shelton and Hedley, 2004), given a long-standing consensus that readers prefer physical paper over screen-based reading  (Tashman and Edwards, 2011a; Sellen and Harper, 2002).
However, none of the existing systems have achieved an on-demand document enhancement. Instead, documents required pre-processing with content manually prepared before being embedded in AR. For example, HoloDoc (Li et al., 2019), PaperTrail (Rajaram and Nebeling, 2022), and Affinity Lens (Subramonyam et al., 2019) did not extract text directly from the given document; instead, content was prepared in advance and loaded based on attached fiducial markers. As a result, these systems are not fully adaptable and deployable to real-world scenarios. Some preliminary works have explored on-demand text analysis using machine learning, such as Dually Noted (Qian et al., 2022) analyzing document structure in real-time, Augmented Math (Chulpongsatorn et al., 2023) extracting graphs and math equations, and SOCRAR (Strecker et al., 2022) using OCR to extract key information. However, none have yet achieved comprehensive and general-purpose document enhancement.
@@
Outside of the context of mixed reality, a wide range of projects and systems have explored reading support tools for screen-based interfaces. Examples including LiquidText (Tashman and Edwards, 2011b), texSketch (Subramonyam et al., 2020), ScholarPhi (Head et al., 2021), Chameleon (Masson et al., 2020), and Marvista (Chen et al., 2022) developed on-screen reading support tools. Similar to our work, Marvista (Chen et al., 2022) helps users read online documents by providing reading aids including summaries, contextual prompt questions, and reading metrics for active reading. Other works studied active and close reading support like Metatation (Mehta et al., 2017) for closed reading, XLibris (Price et al., 1998) with free form annotation, Matulic and Norrie (Matulic and Norrie, 2012) to support navigation, and GatherReader (Hinckley et al., 2012) with information gathering. Other works also support literature review and reference search for science papers (Palani et al., 2023; Kang et al., 2023; Chang et al., 2023). Lastly, works like Explorable Explanations (Victor, 2011) and Potluck (Litt et al., 2022) try to enhance static documents with interactive content to improve reading comprehension through exploration. While these past works are limited to screen-based interfaces, this paper explores the augmented reading design for AR-based interfaces.
@@ 1.2. Content Summarization
In the field of natural language processing, several works have employed automatic summarization to support readers (El-Kassas et al., 2021; Zhang et al., 2023). These approaches provide various summarization methods, including abstractive and extractive summaries generated from document photos (Benharrak et al., 2022), real-time summarization of user writing for quick text iterations (Dang et al., 2022), enhancing reading comprehension through automatic summarization (Chen et al., 2022), and responsive text summarization for adapting document summaries to screen sizes ranging from large displays to small watch faces (Leiva, 2018). With the recent advances in large language models like GPT-4, such automatic summarization tasks have become more accurate, robust, and accessible. For example, Goyal et al. (Goyal et al., 2022) showed that GPT-3 summaries were preferred by humans over those generated by fine-tuned models. Various types of summarization have been explored using GPT-3, such as opinion summarization (Bhaskar et al., 2022), news summarization (Goyal et al., 2022), and medical dialogue summarization (Chintagunta et al., 2021) while others have built custom frameworks for feeding long texts into GPT-3 to elicit more finely-controlled summarizations (Payne, 2022).
@@
Researchers have also explored interactive summarization approaches based on context and users’ needs. For example, Hoeve et al. (ter Hoeve et al., 2020) studied Q&A interactions to assist users with a document-centric question-answering paradigm, finding that information-seeking was a common query. Semantic Reactor (Google, [n. d.]), developed by Google, enables users to ask questions about documents in natural language by fine-tuning BERT (Devlin et al., 2019) using the SQuAD dataset (Rajpurkar et al., 2018). VERSE (Vtyurina et al., 2019) is another Q&A system that allows users to ask specific types of questions about a document but is limited by its inability to summarize. Hoeve et al. (ter Hoeve et al., 2020) identified various question types asked by participants regarding documents, with factual/summarization questions being the most frequent. Some question types include document-related, factoid, mechanical (answerable by rule-based methods), factual, yes/no, navigational, and summary questions. Voice assistants like Cortana, Alexa, Siri, VERSE (Vtyurina et al., 2019), Firefox Voice (Cambre et al., 2021), and Pushpak (Holani et al., 2019) offer efficient query modalities, as speech input has been found to be significantly faster than manual input, especially for mixed reality devices (Ruan et al., 2016; Adhikary and Vertanen, 2021). Our goal is to build on top of this prior work, developing and exploring automatic and interactive summarization techniques for on-demand and real-time summaries embedded in documents.
The limitations of prior works such as Holodoc (Li et al., 2019), as explained before, prevented the study of long term and in-the-wild use of MR reading assistants. To our knowledge GazePointAR (Lee et al., 2024) is the closent work which explored long-term use of pronoun disambiguation in MR voice assistants. Still, it is unclear how can LLM powered MR assistants are used in everyday life to assist users in various types of reading and knowledge work tasks. We aim to fill this gap. In this paper, we contribute to the first demonstration of on-demand document enhancement by leveraging large language models. The goal of this paper is not only to demonstrate its potential but also to uncover its limitations and challenges from a user experience perspective through iterative implementation and multiple user evaluations.
@@ 2. Overview of Our Research Approach
To uncover insights into on-demand reading assistants, we employed a three-phase iterative design process to investigate user interactions and experiences with RealitySummary. Our approach consisted of a usability study on the initial prototype (Study 1: RS-Doc), followed by an in-the-wild study on a second prototype (Study 2: RS-Wild), and concluded with an author-conducted diary study using a third prototype (Study 3: RS-Diary). Each phase was designed to build upon the insights gained from the previous round, allowing for continuous refinement of the design and functionality.
We chose this multi-method, iterative approach to capture a comprehensive understanding of user interactions across various contexts. The initial usability study provided controlled insights into the basic functionality and user experience of RealitySummary. The subsequent in-the-wild study allowed us to observe how users interacted with the improved, more reliable prototype in everyday environments, uncovering real-world usage patterns and challenges. Finally, the diary study offered deep, longitudinal insights into the integration of RealitySummary into daily life, allowing for nuanced observations of subtle interaction patterns and evolving user needs. This progression from controlled to naturalistic settings, coupled with iterative prototyping, allowed us to holistically explore the complex dynamics of AR-AI interactions while addressing emerging design considerations and ethical implications at each stage.
3. Study 0: Elicitation Study
Before implementing our system, we conducted a design elicitation study to explore potential designs for an AI-assisted augmented reading assistant. We held formative design workshops with six participants (4 males and 2 females) who had various design backgrounds as undergraduate and graduate HCI researchers. The study involved two tasks, each lasting approximately 15-20 minutes, with the entire session taking about 60 minutes.
3.1. Method
We asked participants to complete a four-part design task where they read and graphically annotated both a document of their own and a Wikipedia article [^1] Source: https://en.wikipedia.org/wiki/Russian_invasion_of_Ukraine. After reading, participants brainstormed ideas for AI-powered augmentations to improve readability. In the first design task, participants annotated a physical printout of their own article, sketching ideas for possible enhancements, and then repeated the task with a printed copy of the Wikipedia article. In the second task, participants were provided with digital copies of the two documents on a virtual whiteboard (Miro) and asked to design a second set of document enhancements, this time with assistance from ChatGPT.
Before the study began, we briefed participants, gave them a reference page of sample visualizations, and demonstrated prompts for ChatGPT that showcased its summarization and enhancement capabilities. Using a think-aloud protocol, the authors took detailed notes and collected all participant sketches. These designs were later collated on a Miro board, where four authors independently categorized and coded them, identifying 20 distinct types of document enhancements based on usage, purpose, popularity, and context. The authors then discussed and refined these categories, re-coding the examples to produce a final set of enhancement types  (Figure 3).
Figure 3. In our elicitation study, participants identified a variety of potential summary and augmentation types based on sketching and ChatGPT exercises using diverse text documents.

### v3

@@ Abstract.
Large Language Models (LLMs) are gaining popularity as reading and summarization aids. However, little is known about their potential benefits when integrated with mixed reality (MR) interfaces to support everyday reading. In this iterative investigation, we developed RealitySummary  [^1] Project Page: https://adigunturu.github.io/RealitySummary_SUI25/, an MR reading assistant that seamlessly integrates LLMs with always-on camera access, OCR-based text extraction, and augmented spatial and visual responses. Developed iteratively, RealitySummary evolved across three versions, each shaped by user feedback and reflective analysis: 1) a preliminary user study to understand reader perceptions (N=12), 2) an in-the-wild deployment to explore real-world usage (N=11), and 3) a diary study to capture insights from real-world work contexts (N=5). Our empirical studies’ findings highlight the unique advantages of combining AI and MR, including always-on implicit assistance, long-term temporal history, minimal context switching, and spatial affordances, demonstrating significant potential for future LLM-MR interfaces beyond traditional screen-based interactions.
@@
journalyear: 2025copyright: ccconference: ACM Symposium on Spatial User Interaction; November 10–11, 2025; Montreal, QC, Canadabooktitle: ACM Symposium on Spatial User Interaction (SUI ’25), November 10–11, 2025, Montreal, QC, Canadadoi: 10.1145/3694907.3765933isbn: 979-8-4007-1259-3/2025/11ccs: Human-centered computing Mixed / augmented reality
Figure 1. RealitySummary is an on-demand mixed reality reading assistant developed over three iterations, with each version extensively shaped by user feedback and reflection. Through this evolution, we highlight key insights and two notable findings on integrating LLMs in MR interfaces: 1) the unique benefits of MR + LLMs, and 2) emergent use cases for MR reading assistants.
@@ 1.1. Reading Augmentation
Human-computer interaction researchers have explored ways to augment physical documents with AR-enhanced content. Examples include HoloDoc (Li et al., 2019), Replicate and Resuse (Gupta et al., 2020), and PaperTrail (Rajaram and Nebeling, 2022), all of which laid the groundwork for augmenting physical paper by superimposing dynamic virtual content onto printed documents. Prior work has demonstrated these concepts on different platforms. For example, Pacer (Liao et al., 2010), MagPad (Xu et al., 2015), Teachable Reality (Monteiro et al., 2023), and Dually Noted (Qian et al., 2022) use mobile AR for text augmentation, whereas Replicate and Resuse (Gupta et al., 2020) and Wiki-TUI (Wu et al., 2007) leverage head-mounted displays to overlay information. Other examples such as DigitalDesk (Wellner, 1991), DocuDesk (Everitt et al., 2008), Matulic et al.’s pen and touch systems (Matulic and Norrie, 2013; Matulic et al., 2013), and QOOK (Zhao et al., 2014) use projection mapping for document augmentation. Finally, cAR (Hincapié-Ramos et al., 2014) leveraged transparent displays for document augmentation to support active reading. At their core, such augmented reading research (Kirner et al., 2012; Billinghurst et al., 2001; Hidayat et al., 2020; Grasset et al., 2008) aims to improve cognitive load (Cheng, 2017), knowledge accumulation (Dünser et al., 2012; Zhang et al., 2019), and spatial visualization (Shelton and Hedley, 2004), given a long-standing consensus that readers prefer physical paper over screen-based reading  (Tashman and Edwards, 2011a; Sellen and Harper, 2002).
However, none of the existing systems have achieved on-demand document enhancement. Here, we define on-demand as the ability to dynamically and interactively enhance documents without requiring prior manual preparation or pre-processing. For example, prior work like HoloDoc (Li et al., 2019), PaperTrail (Rajaram and Nebeling, 2022), and Affinity Lens (Subramonyam et al., 2019) did not extract text directly from the given document. Instead, content was prepared in advance and loaded based on attached fiducial markers. As a result, these systems are not fully adaptable and deployable to real-world scenarios. Some preliminary work has explored on-demand text analysis using machine learning, such as Dually Noted (Qian et al., 2022) analyzing document structure in real-time, Augmented Math (Chulpongsatorn et al., 2023) and Augmented Physics (Gunturu et al., 2024) extracting diagrams and math equations, and SOCRAR (Strecker et al., 2022) using OCR to extract key information. However, none have yet achieved comprehensive and general-purpose document enhancement.
@@
Outside of the context of mixed reality, a wide range of projects and systems have explored reading support via screen-based interfaces. Examples include LiquidText (Tashman and Edwards, 2011b), texSketch (Subramonyam et al., 2020), ScholarPhi (Head et al., 2021), Chameleon (Masson et al., 2020), and Marvista (Chen et al., 2022). Similar to our work, Marvista (Chen et al., 2022) helps users read online documents by providing reading aids including summaries, contextual prompt questions, and reading metrics. Other research has studied active and close reading support, including Metatation (Mehta et al., 2017) for close reading, XLibris (Price et al., 1998) for free form annotation, Matulic and Norrie’s tabletop interfaces (Matulic and Norrie, 2012) to support navigation, and GatherReader (Hinckley et al., 2012) for information gathering. Other systems have supported literature review and reference search for scientific papers (Palani et al., 2023; Kang et al., 2023; Chang et al., 2023). Lastly, works like Explorable Explanations (Victor, 2011) and Potluck (Litt et al., 2022) try to enhance static documents with interactive content to improve content comprehension through exploration. While this past work is limited to screen-based interfaces, our paper explores the augmented reading design for AR-based interfaces.
@@ 1.2. Content Summarization
In the field of natural language processing, a variety of tools have employed automatic summarization to support readers (El-Kassas et al., 2021; Zhang et al., 2023). These approaches provide various summarization methods, including abstractive and extractive summaries generated from document photos (Benharrak et al., 2022), real-time summarization of user writing for quick text iterations (Dang et al., 2022), enhanced reading comprehension through automatic summarization (Chen et al., 2022), and responsive text summaries for adapting documents to screen sizes ranging from large displays to small watch faces (Leiva, 2018). With recent advances in large language models like GPT-4 and GPT-5, such automatic summarization tasks have become more accurate, robust, and accessible. For example, Goyal et al. (Goyal et al., 2022) showed that GPT-3 summaries were preferred by humans over those generated by fine-tuned models. Various types of summarization have been explored using GPT-3, such as opinion summarization (Bhaskar et al., 2022), news summarization (Goyal et al., 2022), and medical dialogue summarization (Chintagunta et al., 2021) while others have built custom frameworks for feeding long texts into GPT-3 to elicit more finely-controlled summarizations (Payne, 2022).
@@
Researchers have also explored interactive summarization approaches based on context and users’ needs. For example, Semantic Reactor (Google, [n. d.]), developed by Google, enables users to ask questions about documents in natural language by fine-tuning BERT (Devlin et al., 2019) using the SQuAD dataset (Rajpurkar et al., 2018). VERSE (Vtyurina et al., 2019) is another Q&A system that allows users to ask specific types of questions about a document but is limited by its inability to summarize. Hoeve et al. (ter Hoeve et al., 2020), meanwhile, identified a diverse questions asked by participants while reading documents, with factual/summarization questions being the most frequent. Other types include document-related, factoid, mechanical (answerable by rule-based methods), factual, yes/no, navigational, and summary questions. Voice assistants like Cortana, Alexa, Siri, VERSE (Vtyurina et al., 2019), Firefox Voice (Cambre et al., 2021), and Pushpak (Holani et al., 2019) offer efficient query modalities, as speech input has been found to be faster than manual input, especially for mixed reality devices (Ruan et al., 2016; Adhikary and Vertanen, 2021).
Reality-Based Information Retrieval
Other recent research has examined visual reality-based information queries (Büschel et al., 2018). For example, GazePointAR (Lee et al., 2024) explored long term use of pronoun disambiguation in visual question answering in mixed-reality. G-Voila (Wang et al., 2024) explored gaze as a medium for question answering in everyday scenarios. While prior work has either focused on document augmentation with manual preparation (Li et al., 2019) or on general mixed-reality LLM-based QnA (Lee et al., 2024), our goal is to build on this foundation by developing and exploring on-demand and real-time reading augmentation within everyday Mixed-Reality settings.
To our knowledge, RealitySummary is the first LLM powered MR reading system to be studied in the wild over an extended period, enabling the collection of rich, longitudinal insights into real-world reading behaviors. This long-term, situated deployment serves as the foundation for our contribution, upon which we highlight implications of on-demand, context-aware mixed reality augmentations.
@@ 2. Overview of Our Research Approach
To uncover insights into on-demand reading assistants, we employed a three-phase iterative design process to investigate reader interactions and experiences with MR reading assistants. These explorations included a usability study using the initial prototype (Study 1: RS-Doc), followed by an in-the-wild study with a second prototype (Study 2: RS-Wild), and an author-conducted diary study with a third prototype (Study 3: RS-Diary). Each phase was designed to build upon the insights gained from the previous round, allowing for continuous refinement of the design and functionality. Insights from each study motivated the design of the subsequent system, and the three studies are presented together in this paper as part of a single continuous investigation, providing a complete picture of the iterative design process and cross-study findings. All studies were conducted under the approval of the University of Calgary ethics board (IRB Approval Number: REB21-2065). Participants provided informed consent prior to participation and could withdraw at any time without penalty.
3. RS-Doc: Design Exploration and Usability Study with the First Prototype
We implemented the first version of our system (RS-Doc), which focused on in-place summarization for physical and digital documents, on a Microsoft HoloLens 2. In this section, we present initial observations from our experiences developing and evaluating RS-Doc as well as new design opportunities highlighted by those explorations.
3.1. Implementation
System Design
RS-Doc captures documents using the HoloLens 2’s internal camera and extracts text with Google Cloud OCR. The extracted text is processed using GPT-4 via OpenAI’s API to identify key information, while spaCy’s named entity recognition (NER) module assists in identifying important entities. The system employs Vuforia for markerless image tracking, and the Mixed Reality Toolkit 3 (MRTK3) SDK for content rendering. Readers interact with the system through voice commands processed via Google’s speech-to-text API, allowing them to modify or query visualizations. The system captures images periodically for OCR, with an option for manual activation, and tracks documents by generating dynamic image targets, ensuring accurate document tracking and augmented content rendering.
Enhancement Types
To design and implement our system, we identified five key categories of AI-generated content enhancements: summarize, compare, augment, extract, and navigate—based on an early elicitation study (see appendix) as well as prior work, such as HoloDoc (Li et al., 2019), PaperTrail (Rajaram and Nebeling, 2022), DuallyNoted (Qian et al., 2022), QOOK (Zhao et al., 2014), and Digital Desk (Li et al., 2019). These categories informed both the conceptual framework and the system features. Summarize condenses text for quick reference, ranging from broad overviews to personalized summaries. This is implemented by sending the current page’s text to GPT-4 for concise summaries, limited to 150 words. Compare transforms data into visual formats by parsing documents. Our system provides the compare feature by generating comparison tables for documents discussing multiple related themes, entities, or concepts, as well as timelines for documents describing sequences of events. Augment enhances content with external data. For instance, HoloDoc (Li et al., 2019) and PaperTrail (Rajaram and Nebeling, 2022) augment papers with external content, such as references, videos, and figures. Similarly, our system augments documents by providing information cards about people and places using Google Search, Image APIs, and Mapbox, based on keywords identified by spaCy. Extract enables readers to pull and store elements like keywords and quotes. For example, DigitalDesk (Wellner, 1991) allows users to extract content from documents for copy-and-paste functionality. RS-Doc implements this feature by generating keyword lists that highlight the most central named entities extracted. Finally, Navigate improves document navigation through features like a table of contents or bookmarks. For instance, QOOK (Zhao et al., 2014) allows users to add bookmarks to specific sections. RS-Doc supports navigation by generating a table of contents, created by the LLM based on the document.
Interactions
After launching RS-Doc, a reader can trigger summaries for an in-view document by saying “give me summaries”. The system then captures an image of the page currently in view, creates a dynamic image target, and generates summary cards of all types relevant to that page. Summaries are displayed around the page using a fixed layout (for example, timeline summaries always appear at the top right of the page and information cards always appear to the left) and remain locked to the page if either the document or reader moves. For example, if a user is reading a magazine about aerial drones, the system automatically detects important keywords such as people, places, and phrases from the page. The system then generates an image, a comparison table, a summary, and information about key people and keywords. Readers can also show or hide specific cards using voice commands (“hide keywords”, “show timeline”, etc.). Note that the system only supports one card of each type at a time. The user can query for a card multiple times, which will re-prompt the LLM (each card type has its own fixed prompt which the user cannot change).
The decision to lock content to the document was motivated by the limitations noted in prior work like Holodoc (Li et al., 2019), where the participants expressed frustration and preferred the content to follow their views. Additionally, prior research, such as Dually Noted (Qian et al., 2022), indicates that the participants preferred not to obstruct the view of the page itself. Our system displays the cards around the page and tracks them in MR with its physical movement.
Document Tracking and Content Generation Performance
To characterize the technical performance of our prototype, we assessed document tracking reliability and system latency. To test tracking, we collected 60 seconds of tracking data each for 20 diverse documents (10 physical, 10 on-screen) including posters, newspapers, patents, and books—then recording what fraction of time the document was correctly registered. Tracking was largely reliable for documents with visuals (M=92%, SD=5.4%) but less so for text-only documents (M=64%, SD=3.7%) which had fewer distinct features. Document tracking was more reliable when visuals (e.g., diagrams or figures) were present, but less reliable for text-only documents. Tracking performed poorly as compared to text detection because Vuforia’s image tracking depends on detecting distinctive feature points, which are more abundant in images and figures than in uniform text regions. Text-only pages often contain repetitive patterns and large blank areas harder to capture with the HoloLens camera that provide fewer robust keypoints, reducing tracking stability. Performance varied with lighting (better in well-lit conditions), target size (larger images performed better), and motion (sudden movements caused temporary tracking loss).
We also measured the average latency of three components: AR rendering, network communication, and API responses. AR rendering averaged 100 ms on HoloLens, while network communication via socket.io averaged 40 ms, indicating responsive performance. API callbacks included GPT-4 requests (2.4 s) and our NLP module using spaCy (670 ms). Since we limited responses to a short response (150 words for a summary), GPT-4 latency remains relatively low. These latencies remained largely consistent across our two later prototypes.
3.2. Study Method
Using the RS-Doc prototype, we conducted an initial usability study involving 12 participants (P1-P12, 6 male and 6 female, aged 19-35) recruited via snowball sampling and university-wide advertisements. We compensated participants \$15 CAD for the one-hour session. All participants reported to be proficient in English and accustomed to reading academic or professional materials. As all were enrolled university students, they met our inclusion criterion of being able to read documents of varied comprehension levels. No additional exclusion criteria were applied. The study aimed to evaluate the system’s usability, explore participants’ preferences for different document enhancement strategies, and identify limitations for future iterations, with real-world use in mind. We conducted the study in a lab environment, where we first introduced participants to the HoloLens 2 and then provided an overview of the system and the set of document enhancements. Participants used our system to explore a variety of materials and media: 1) an A4-sized magazine and an A3-sized book in physical format, and 2) a digital document of their choice. The physical magazine and book contained the same reading content for all participants, while the digital documents were selected by participants to reflect their diverse preferences and included blog posts, course assignments, news articles, Wikipedia entries, and digital books. The physical magazine and book each contained fixed content, and the same materials were used across all participants. This ensured that differences in responses could be compared consistently within each medium, without variability introduced by differing texts. This study aimed to investigate participants’ interactions and preferences across different reading materials and formats, rather than focusing on standardized tasks; therefore, we did not explicitly measure or control for difficulty levels. We tasked the participants with reading and understanding the documents provided to them, using RealitySummary in whichever way they saw fit and using whatever combination of enhancements best suited the document.
Throughout the study, we encouraged participants to think aloud and documented their experiences. Following the tasks, we also conducted semi-structured interviews and a Likert scale questionnaire, including a System Usability Scale (SUS) (Brooke et al., 1996). We then performed a reflexive thematic analysis (Clarke and Braun, 2014) using transcripts of both the think-aloud sessions and interviews, from which we identified and refined key themes.